OpenStudy (anonymous):

Solve for x. √8(√2-x)=11

OpenStudy (anonymous):

\[\sqrt{8} *\sqrt{2} = \sqrt{16} = 4\] \[4 - \sqrt{8}x = 11\] \[-\sqrt{8}x = 7\] No pretty answer I'm afraid